OpenStudy (anonymous):

Cartesian and polar coordinates are related through the transformation equations {x = r cos Theta or {r^2= x^2 + y^2 {Y = r sin theta { tan theta = y/x A. Evaluate the partial derivatives Xr, Yr,Xtheta, and Ytheta. B. Evaluate the partial derivatives Rx, Ry, Theta x, Theta y. C. For a function z = f(x,y), find z, and Ztheta, where x and y are expressed in terms of r and theta. D. for a function z = g(r,theta), find Zx, and Zy, where r and theta are expressed in terms of x and y. E. show that (∂z/∂x)^2 + (∂z/∂y)^2 = (∂z/∂r)^2 + 1/r^2 (∂z/∂theta)^2

OpenStudy (amistre64):

polars give me a headache still :)

OpenStudy (anonymous):

we haven't learned this and she gives it as a quiz problem smh

OpenStudy (amistre64):

a and b are simple enough id think

OpenStudy (amistre64):

z is a circle

OpenStudy (anonymous):

lol i am not that smart. still trying to figure out the answer for other one

OpenStudy (amistre64):

partials mean that the only thing changing is variable we are diffing with. wrt to r means that every thing that aint r is not changing, when things dont change they are consant. Xr = cos(t) ; Yr = sin(t)

OpenStudy (amistre64):

x = r cos(t) ; Xr means that only r changes, everything else remains consant Xr = cos(t) then

OpenStudy (amistre64):

Xt would mean that r doesnt change and remains constant .... Xt = -r sin(t)

OpenStudy (anonymous):

this is for a and b right

OpenStudy (amistre64):

that is the concepts behind a and b yes

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

z = f(x,y); this ones a bit vague to me since it really doesnt provide us with a structure for f(x,y) other than to assume we use x and y in the rule for the function

OpenStudy (amistre64):

z = xy - x^2 +3y would be a function of z in x and y

OpenStudy (amistre64):

if we allow z to mimic the results of the vector function: z =<rcos(t),rsin(t)> we get a circle centered at the origin of radius r

OpenStudy (anonymous):

i am so confused

OpenStudy (amistre64):

to which: z^2 = r^2 -(x/r)^2 -(y/r)^2

OpenStudy (amistre64):

lol, im just a little confused ;)

OpenStudy (anonymous):

lol i have no clue whats going on

OpenStudy (amistre64):

yeah, after a and b im lost :) srry

OpenStudy (anonymous):

i am even lost on a and b